OpenStudy (anonymous):

Zoe looked at a map of a small island where her cousin lives. The map indicated that 1 inch = 5 miles. On the map the distance between her cousins house and the nearest beach is 2 1/3 inches. What is the distance between her cousins house and the nearest beach? 11 miles 11 2/3 7 1/3 12 1/3 Please explain I will give medal and be ur fan!!

OpenStudy (anonymous):

??? any1 please ???

OpenStudy (jagatuba):

1 1/3 * 5

OpenStudy (anonymous):

Like 13 miles, I believe.

OpenStudy (jagatuba):

*2 1/3

OpenStudy (anonymous):

and how do I do dat...?

OpenStudy (anonymous):

im not so great at math....

OpenStudy (jagatuba):

\[\frac{ 7 }{ 3 }\times \frac{ 15 }{ 3 }\]

OpenStudy (anonymous):

oooh ok :)

OpenStudy (anonymous):